222 R 422 W 412 M 6 V 402 2V = a) Determine the value of R for maximum power transfer to R. b) Find the maximum power to R.
Answer #1

first of all I would like to eseplain about The vinens Theorem and Maximum Power Transfer Theorem and with the help of thatThevenios Theorem at states that lineaa bilateral network any 00 in series with the load (R) as follows: aplaced by a singleMaximum Power Transfer Theorem: It states that for a thevenin Gruivalent circuit, the maximum power R₂ =R Th transfers to theGixen circuit diagram R 7 42 OV 4.12 2V Here I am considering R? load asistance as R. -R To find Vith! follows: After RemovinBy applying Kch at ve node: V Ve-V6 Ve-Va 2 4- Vb 4 ****** [3] By substituting (No=2V V[1] + Vo[] X (VG = 8 4 Vc + va(-1/2) =By substituting ve cin eruation o (16 (34) -3)+ V.(-) = 1/2 3.5V Va 35V By substituting Va cin ve eqration: V = (35) (372) 5)To find Rini 1 RTLY 3 1 42 22 - 2 42 1 42 - TH I Here I an dividing diagram dividing diagram in to 3 parts for calculation paPast 0 42 Thi Equivalent Resistance of this part CR.). RA 411141 - + Ri to 2/4=14 Ri > Rican 202 7 Ri= goon w Part 22 42 By s• Ra -> M Part 3 R Th 2.12 M - ny 42 2 TIT By substituting part a with its equivalent Resistance RTU 42 ARS 9 Th Ill Rth = 4.Maximuern power from transfer theorem We know that maximum power transfer to load (R = ) R should be earual to Rth R = Rth =2